# Multimodal classification of forest biodiversity potential from 2D orthophotos and 3D airborne laser scanning point clouds

### v2

@@ Abstract
Assessment of forest biodiversity is crucial for ecosystem management and conservation. While traditional field surveys provide high-quality assessments, they are labor-intensive and spatially limited. This study investigates whether deep learning-based fusion of close-range sensing data from 2D orthophotos and 3D airborne laser scanning (ALS) point clouds can reliable assess the biodiversity potential of forests. We introduce the BioVista dataset, comprising 44 3784437844\,37844 378 paired samples of orthophotos and ALS point clouds from temperate forests in Denmark, designed to explore multimodal fusion approaches. Using deep neural networks (ResNet for orthophotos and PointVector for ALS point clouds), we investigate each data modality’s ability to assess forest biodiversity potential, achieving overall accuracies of 76.7% and 75.8%, respectively. We explore various 2D and 3D fusion approaches: confidence-based ensembling, feature-level concatenation, and end-to-end training, achieving overall accuracies of 80.5%, 81.4% and 80.4% respectively. Our results demonstrate that spectral information from orthophotos and structural information from ALS point clouds effectively complement each other in forest biodiversity assessment.
@@ 1.1 Background and motivation
Forest biodiversity is fundamental to ecosystem functioning, supporting critical processes such as nutrient cycling, carbon sequestration, water regulation, and soil formation, which are vital for planetary health and human well-being (Goldstein et al., 2020; Thompson et al., 2009). This essential biodiversity is facing unprecedented threats globally due to accelerating human activities, including deforestation, habitat fragmentation, and climate change (Aerts and Honnay, 2011; FAO and UNEP, 2020).
The severity of this ongoing loss has spurred significant regional and international initiatives, like the EU Biodiversity Strategy for 2030 (European Commission, 2020) and the UN’s Global Forests Goals and Sustainable Development Goals (SDGs) (United Nations Department of Economic and Social Affairs, 2021, 2017), aimed at increasing protection and enhancing biodiversity. Effectively prioritizing conservation efforts and resources under these initiatives requires scalable and accurate methods for identifying areas of high biodiversity value (Kerry et al., 2022), a challenge this study seeks to address.
@@ 1.2 Traditional assessment methods and limitations
Traditional forest biodiversity assessment methods rely heavily on resource-intensive manual fieldwork. Established approaches like the Forest Health Monitoring Program in the United States (Potter and Conkling, 2022), the RAINFOR Amazon Forest Inventory Network methodology in South America (Malhi et al., 2002) and the International Co-operative Program on Assessment and Monitoring of Air Pollution Effects on Forests in Europe (Michel et al., 2022) provide valuable qualitative insights through detailed ground surveys. However, these methods are limited in their spatial and temporal coverage due to the intensive nature of manual tree measurement, plant identification, and habitat assessment protocols. As a result, large-scale biodiversity assessments often rely on extrapolation from sparse sample plots (Kangas et al., 2018; Nord-Larsen et al., 2023).
@@ 1.3 Remote sensing and deep learning approaches
Remote sensing and close-range remote sensing data, such as multispectral images and LiDAR point clouds from airborne sensors, combined with recent advancements in deep learning methods for processing 2D imagery (He et al., 2015; Dosovitskiy et al., 2021; Woo et al., 2023) and 3D point clouds (Qi et al., 2017; Guo et al., 2021; Deng et al., 2023) offers promising capabilities for rapid and large-scale forest assessment (Zhu et al., 2017; Yuan et al., 2020; Brandt et al., 2024).
@@
Recent deep learning approaches have shown to improve performance in tree species classification (Bjerreskov et al., 2021) as well as other forest-related variables, such as tree counting (Cheng et al., 2024), tree height prediction (Li et al., 2023), forest canopy cover estimation (Liu et al., 2023; Reiner et al., 2023), tree mortality estimation (Khatri-Chhetri et al., 2024), land cover classification (Kussul et al., 2017) directly from various remote sensing imagery.
@@
Similarly, deep learning-based classification of tree species classification (Xi et al., 2020) and assessment of forest structural characteristics like aboveground biomass (Oehmcke et al., 2024) and canopy height estimation (Lang et al., 2023), directly from LiDAR point clouds, has seen recent and promising advancements as well.
@@
Fusing complementary 2D imagery and 3D LiDAR data has been shown to improve results over using either modality alone for e.g. land cover classification and above-ground biomass estimation  (Jia, 2015; Rastiveis, 2015; Hang et al., 2020; Daneshtalab et al., 2019; Zhang et al., 2019). But common fusion strategies exhibit notable limitations. One widespread approach involves extracting hand-crafted geometric or statistical features from the 3D point clouds to combine with the 2D image data, frequently using traditional machine learning classifiers (Jia, 2015; Rastiveis, 2015). This manual feature engineering requires domain expertise and can be suboptimal as it risks discarding potentially valuable patterns present in the raw point clouds. Other methods convert or project the 3D LiDAR data into 2D representations (e.g., sparse images, normalized Digital Surface Models, intensity layers) to facilitate fusion, often enabling the use of standard image-based CNNs (Hang et al., 2020; Daneshtalab et al., 2019). However, this strategy inevitably sacrifices the rich 3D structural and spatial relationships inherent in the original point cloud data by avoiding direct processing in its native format.
@@
While many forest-related variables have been estimated or assessed, we have not identified studies in the literature that directly tackle forest biodiversity assessment from 2D imagery, 3D point clouds, or their combination using deep learning or other machine learning techniques.
@@ 1.4 Research objectives and contributions
Forest biodiversity is an umbrella term which refers to the variety of living organisms such as trees, plants, animals, and microbes found in forest environments (Kumar, 2022). In this work, we define forest biodiversity in terms of presence of 11 High Nature Value (HNV) proxy features, introduced by (Johannsen et al., 2015), which have been statistically shown to have a high correlation with habitats of endangered species (animals, plants and fungi) (Ejrnæs et al., 2014) within Danish forests. We use the term forest biodiversity potential, since our forest biodiversity definition is based on presences of proxy features, which are indicators of biodiversity potential, rather than being a direct inventory of biodiversity. Further details on the HNV proxy features are provided in Section 2.2 and the list of the 11 features are presented in Table 1.
@@
    The BioVista dataset: Development the BioVista dataset, containing 44 3784437844\,37844 378 pairs of 2D orthophotos and 3D ALS point clouds, labeled as high- or low forest biodiversity potential. This dataset and the associated annotations are made publicly available and provide a valuable resource for close-range remote sensing-based forest biodiversity assessment.
@@
    Multimodal fusion: We evaluate deep learning-based multimodal fusion methods that combine 2D spectral- and 3D structural data and clearly demonstrate how the two data modalities complement each other effectively for forest biodiversity assessment.
@@
The remainder of this paper is organized as follows: Section 2 presents the BioVista dataset, including data collection, preprocessing and preparation. Section 3 details our deep learning methodology, covering both single-modality approaches (2D orthophoto and 3D ALS point cloud classification) and multimodal fusion strategies. Section 4 presents our experimental results and performance analysis. Section 5 discusses limitations and potential improvements of our approach and Section 6 concludes the work with key findings and future research directions.
@@ 2 Dataset
Figure 2: The BioVista dataset. (a) Geographic distribution of forest areas used for sampling training, validation and test sample plots. (b) Example of two forest areas which contain 84 and 39 circular samples. Samples have a diameter of 30 meters and samples from the same forest area belong to the same class. (c) A single sample is represented as a 2D orthophoto, where 1 pixel is equal to 12,5 cm on the ground, and a 3D ALS point cloud with 8 points/m².
@@ 2.1 The BioVista dataset
The BioVista dataset, introduced in this study, combines High Nature Value (HNV) proxies (Johannsen et al., 2015), described in detail in Section 2.2, with 2D orthophotos and 3D ALS point clouds to create a comprehensive resource for forest biodiversity potential classification. The dataset is a unique, high-quality resource for forest biodiversity assessment, made possible by Denmark’s tradition of comprehensive forest registries (Nord-Larsen et al., 2023), extensive geographic data collection, and open data policies (Hansen et al., 2011).
The dataset comprises 44 3784437844\,37844 378 pairs of 2D orthophoto images and 3D ALS point clouds derived from forests in Denmark, labeled as either high- or low biodiversity potential. An overview of the dataset, geographic location of forest areas and sampling methodology can be seen in Figure 2. This shape and size of samples conforms to the test plot standard used in e.g., the National Forest Inventory (NFI) of Denmark (Alban et al., 2019), and is comparable to other nordic NFI standards (Kangas et al., 2018). One sample covers an area of $\sim 707m^{2}$ and the rational is to have a large enough area to include a diverse selection of both small and larger trees and vegetation within a single sample, as well as variations in soil, ground cover, and other ecological factors within the same area.
2.2 Biodiversity ground truths
Figure 3: HNV forest distribution. (a) High Nature Value (HNV) forest map of Denmark (Johannsen et al., 2015), which shows the spatial distribution of forest areas with HNV scores from 1 to 10 where 1 is worst and 10 is best. (b) Quantitative distribution showing that forest areas with HNV scores from 1-3 comprise nearly 75% of Danish forests, while areas with scores from 7-10 represent approx. 2%.
We use the HNV forest map (Johannsen et al., 2015), seen in Figure 3, as the foundation for gathering ground truth labels for our samples. The HNV forest map is based on 11 features that serve as proxies for forest biodiversity, combining historical and current data. Proxy features include among others canopy height variation, coastal proximity, tree- and plant richness, canopy cover, occurrence of large trees, and occurrence of inner forest edges. A short description of each proxy feature is shown in Table 1, but we refer to (Johannsen et al., 2015) for exhaustive descriptions and details.
@@ 2.2 Biodiversity ground truths
Height Variation Variation in tree height/age within 25$\times$25m plots (LiDAR based), indicating structural diversity and more habitats.
Woody Plant Richness Variety of woody plant species in 4$\times$4 km cells (NFI based), suggesting diverse conditions and less intensive forestry.
Large Trees Presence of trees $>60$ cm diameter at breast height in 4$\times$4 km cells (NFI based), indicating continuity and providing specific habitats.
@@
Nature Share 40% Landscape context where at least 40% of the surrounding 1$\times$1 km area is natural habitat, important for species connectivity.
Nature Share 80% Landscape context where at least 80% of the surrounding 1$\times$1 km area is natural habitat, indicating very high connectivity.
@@
Table 1: HNV forest proxy features. A short summary description of each of the 11 proxy features of the HNV forest map. we refer to (Johannsen et al., 2015) for an exhaustive description.
We define low forest biodiversity potential as areas with HNV scores from 1–3 and high forest biodiversity potential as areas with an HNV scores from 7–10, while ignoring forest with HNV scores between 4 and 6. This approach allows for a more focused comparison of deep learning-based classification methods trained on 2D orthophotos, 3D ALS point clouds, and a combination of the two, while still evaluating the forest biodiversity assessment capabilities of deep learning-based methods, on a coarse-grained scaled.
@@
The 2D orthophotos used in this project are derived from yearly nationwide aerial photography of Denmark and are part of GeoDanmark’s geographic administration framework (GeoDanmark, 2024). Orthophotos are created by adjusting each pixel according to the camera’s position and the elevation model of the Earth’s surface, ensuring that the images are geometrically corrected and can be used for precise measurements.
@@
The 3D ALS point clouds used in this study is part of the Danish Height Model dataset (Klimadatastyrelsen, 2024). This comprehensive dataset covers the entire country of Denmark and is continuously updated through yearly systematic airborne laser scanning. As of 2018 and forward, 1/5 of the country is updated annually. In this research we use ALS point clouds from 2019, 2020, 2021, 2022 and 2023 depending on the area of the forest, as illustrated in Figure 5. For each forest area, we ensure temporal consistency by using orthophotos captured in the same year as the corresponding ALS point cloud data, allowing for fair comparison between the two modalities and ensuring that both data sources represent the forest conditions from the same time period. The point clouds are available in tiles spanning 1$\times$1 km and the point density averages 8 points/m2 for data, with a horizontal accuracy of 0.15 m and a vertical accuracy of 0.05 m.

### v3

@@ Abstract
Assessment of forest biodiversity is crucial for ecosystem management and conservation. While traditional field surveys provide high-quality assessments, they are labor-intensive and spatially limited. This study investigates whether deep learning-based fusion of close-range sensing data from 2D orthophotos and 3D airborne laser scanning (ALS) point clouds can reliable assess the biodiversity potential of forests. We introduce the BioVista dataset, comprising 44 37844\,378 paired samples of orthophotos and ALS point clouds from temperate forests in Denmark, designed to explore multimodal fusion approaches. Using deep neural networks (ResNet for orthophotos and PointVector for ALS point clouds), we investigate each data modality’s ability to assess forest biodiversity potential, achieving overall accuracies of 76.7% and 75.8%, respectively. We explore various 2D and 3D fusion approaches: confidence-based ensembling, feature-level concatenation, and end-to-end training, with the latter achieving an overall accuracies of 82.0% when separating low- and high potential forest areas. Our results demonstrate that spectral information from orthophotos and structural information from ALS point clouds effectively complement each other in the assessment of forest biodiversity potential.
@@ 1.1 Background and motivation
Forest biodiversity is fundamental to ecosystem functioning, supporting critical processes such as nutrient cycling, carbon sequestration, water regulation, and soil formation, which are vital for planetary health and human well-being (goldstein2020protecting; thompson2009forest). This essential biodiversity is facing unprecedented threats globally due to accelerating human activities, including deforestation, habitat fragmentation, and climate change (aerts2011forest; fao-state-of-worlds-forests-2020).
The severity of this ongoing loss has spurred significant regional and international initiatives, like the EU Biodiversity Strategy for 2030 (eu-biodiversity-strategy-2020) and the UN’s Global Forests Goals and Sustainable Development Goals (SDGs) (global-forest-goals-2011; sustainable-development-goals-2017), aimed at increasing protection and enhancing biodiversity. Effectively prioritizing conservation efforts and resources under these initiatives requires scalable and accurate methods for identifying areas of high biodiversity value (kerry-overview-remote-monitoring-biodiversity-conservation-2022), a challenge this study seeks to address.
@@ 1.2 Traditional assessment methods and limitations
In general, biodiversity is difficult to quantify, and we often have to resort to measurements of proxy variables that indicate biodiversity potential. Traditional forest biodiversity assessment methods rely heavily on resource-intensive manual fieldwork. Established approaches like the Forest Health Monitoring Program in the United States (usda-forest-health-monitoring-2022), the RAINFOR Amazon Forest Inventory Network methodology in South America (malhi-rainfor-2002) and the International Co-operative Program on Assessment and Monitoring of Air Pollution Effects on Forests in Europe (michel-icp-forest-2022) provide valuable qualitative insights through detailed ground surveys. However, these methods are limited in their spatial and temporal coverage due to the intensive nature of manual tree measurement, plant identification, and habitat assessment protocols. As a result, large-scale biodiversity assessments often rely on extrapolation from sparse sample plots (kangas-forest-inventories-in-nordic-countris-2018; nord-larsen-nfi-2022).
@@ 1.3 Remote sensing and deep learning approaches
Remote sensing and close-range remote sensing data, such as multispectral images and LiDAR point clouds from airborne sensors (ALS), combined with recent advancements in deep learning methods for processing 2D imagery (kaiming-resnet; dosovitskiy-2021-vistion-transformer; Woo-2023-ConvNeXtV2) and 3D point clouds (qi-pointnet++; guo-2021-deep-learning-for-3d-point-clouds; deng-pointvector) offers promising capabilities for rapid and large-scale forest assessment (zhu-2017-deep-learning-in-remote-sensing; yuan2020deep; brandt:24b).
@@
Recent deep learning approaches have shown to improve performance in tree species classification (bjerreskov-2021-tree-species) as well as other forest-related variables, such as tree counting (cheng2024treecounting), tree height prediction (li2023treeheightprediction), forest canopy cover estimation (liu2023treecover; reiner2023treecover), tree mortality estimation (2024-Khatri-Chhetri-tree-mortality), land cover classification (kussul2017-landcover-and-crop-classification) directly from various remote sensing imagery.
@@
Similarly, deep learning-based classification of tree species (xi-2020-tree-species) and assessment of forest structural characteristics like aboveground biomass (oehmcke2024biomassregression) and canopy height (lang-2023-canopy-height), directly from LiDAR point clouds, has seen recent and promising advancements as well.
@@
Fusing 2D imagery and 3D LiDAR data can improve prediction results compared to using the modalities alone. This has, for example, been shown for land cover classification and above-ground biomass estimation (2015-jia-land-classification; rastiveis-2015-lidar-aerial-fusion-urban-classification; hang-2020-lidar-hyperspectral-fusion; daneshtalab-2019-cnn-fusion-aerial-lidar; zhang2019deep). But common fusion strategies exhibit notable limitations. One widespread approach involves extracting hand-crafted geometric or statistical features from the 3D point clouds to combine with the 2D image data, frequently using traditional machine learning classifiers (2015-jia-land-classification; rastiveis-2015-lidar-aerial-fusion-urban-classification). Similarly, rs10020338 combined 2D spectral and 3D structural data through feature engineering and traditional machine learning to assess forest biodiversity potential. This manual feature engineering requires domain expertise and can be suboptimal as it risks discarding potentially valuable patterns present in the raw point clouds and images. Other methods convert or project the 3D LiDAR data into 2D representations (e.g., sparse images, normalized Digital Surface Models, intensity layers) to facilitate fusion, often enabling the use of standard image-based CNNs (hang-2020-lidar-hyperspectral-fusion; daneshtalab-2019-cnn-fusion-aerial-lidar). However, this strategy inevitably sacrifices the rich 3D structural and spatial relationships inherent in the original point cloud data by avoiding direct processing in its native format.
@@
While many forest-related variables have been estimated or assessed, we have not identified studies in the literature that tackle the assessment of forest biodiversity potential directly from 2D imagery, 3D point clouds, or their combination using deep learning or other machine learning techniques.
@@ 1.4 Research objectives and contributions
Forest biodiversity is an umbrella term which refers to the variety of living organisms such as trees, plants, animals, and microbes found in forest environments (Kumar2022EcologicalRelevance). In this work, we define forest biodiversity in terms of presence of 11 High Nature Value (HNV) proxy features, introduced by (johannsen-hnv-skov-2015), which have been statistically shown to have a high correlation with habitats of endangered species (animals, plants and fungi) (ejrnaes-2014-biodiversitetskort) within Danish forests. Endangered species are great indicators of healthy ecosystems and biodiversity as their presence signifies a stable, resilient, and high-quality ecosystem capable of supporting not only them but a wide array of other, more common species (ejrnaes-2014-biodiversitetskort). We use the term forest biodiversity potential, since our forest biodiversity definition is based on presences of proxy features, which are indicators of biodiversity potential, rather than being a direct inventory of biodiversity. Further details on the HNV proxy features are provided in Section 2.2 and the list of the 11 features are presented in Table 1.
@@
    The BioVista dataset: The BioVista dataset was developed, containing 44 37844\,378 pairs of 2D orthophotos and 3D ALS point clouds, labeled as high- or low forest biodiversity potential. This dataset and the associated annotations are made publicly available and provide a valuable resource for close-range remote sensing-based assessment of forest biodiversity potential.
@@
    Multimodal fusion: We evaluate deep learning-based multimodal fusion methods that combine 2D spectral- and 3D structural data and clearly demonstrate how the two data modalities complement each other effectively for assessment of forest biodiversity potential.
@@
The remainder of this paper is organized as follows: Section 2 presents the BioVista dataset, including data collection, preprocessing and preparation. Section 3 details our deep learning methodology, covering both single-modality approaches (2D orthophoto and 3D ALS point cloud classification) and multimodal fusion strategies. Section 4 presents our experimental results and performance analysis. Section 5 discusses limitations and potential improvements of our approach before we conclude in Section 6.
@@ 2 Dataset
Figure 2: The BioVista dataset. (a) Geographic distribution of forest areas used for sampling training, validation and test sample plots, shown on the orthophoto map of Denmark. (b) Example of two forest areas which contain 84 and 39 circular samples. Samples have a diameter of 30 meters and samples from the same forest area belong to the same class. (c) A single sample is represented as a 2D orthophoto, where 1 pixel is equal to 12,5 cm on the ground, and a 3D ALS point cloud with 8 points/m².
@@ 2.1 The BioVista dataset
The BioVista dataset, introduced in this study, combines High Nature Value (HNV) proxies (johannsen-hnv-skov-2015), described in detail in Section 2.2, with 2D orthophotos and 3D ALS point clouds to create a comprehensive resource for forest biodiversity potential classification. The dataset is a unique, high-quality resource for forest biodiversity potential, made possible by Denmark’s tradition of comprehensive forest registries (nord-larsen-nfi-2022), extensive geographic data collection, and open data policies (hansen-e-governance-2011).
Forests cover about 636,079 ha of Denmark, which corresponds to a forest fraction of 14.8%, according to bjerreskov-2021-tree-species. The fractions of conifer and broadleaf forests were estimated to be 45% and 55% of the forest area, respectively (bjerreskov-2021-tree-species). The most common broadleaf trees in Denmark are (in order) beech, oak, birch, Sycamore maple, and ash; the most common conifers are Norway spruce, pine species, Sitka spruce, Nordmann fir, and Noble fir (nord-larsen-nfi-2022). Several tree height maps of Denmark have been made available recently, we refer to liu2023treecover and li2023treeheightprediction.
The dataset comprises 44 37844\,378 pairs of 2D orthophoto images and 3D ALS point clouds derived from forests in Denmark, labeled as either high- or low biodiversity potential. An overview of the dataset, geographic location of forest areas and sampling methodology can be seen in Figure 2. This shape and size of samples conforms to the test plot standard used in, e.g., the National Forest Inventory (NFI) of Denmark (alban-nfi-field-instructions), and is comparable to other nordic NFI standards (kangas-forest-inventories-in-nordic-countris-2018). One sample covers an area of $\sim 707$ m2 and the rational is to have a large enough area to include a diverse selection of both small and larger trees and vegetation within a single sample, as well as variations in soil, ground cover, and other ecological factors within the same area.
2.2 Biodiversity potential labels
Figure 3: HNV forest distribution. (a) High Nature Value (HNV) forest map of Denmark (johannsen-hnv-skov-2015), which shows the spatial distribution of forest areas with HNV scores from 1 to 10 where 1 is worst and 10 is best , shown on the orthophoto map of Denmark. (b) Quantitative distribution showing that forest areas with HNV scores from 1-3 comprise nearly 75% of Danish forests, while areas with scores from 7-10 represent approx. 2%.
We use the HNV forest map (johannsen-hnv-skov-2015), seen in Figure 3, as the foundation for gathering ground truth labels for our samples. The HNV forest map is based on 11 features that serve as proxies for forest biodiversity, combining historical and current data. Proxy features include, among others, canopy height variation, coastal proximity, tree and plant richness, forest cover, occurrence of large trees, and occurrence of inner forest edges. A short description of each proxy feature is shown in Table 1, but we refer to johannsen-hnv-skov-2015 for details.
@@ 2.2 Biodiversity potential labels
Height Variation Variation in tree height/age within $\times25$m plots (LiDAR based), indicating structural diversity and more habitats.
Woody Plant Richness Variety of woody plant species in $\times4$ km cells (NFI based), suggesting diverse conditions and less intensive forestry.
Large Trees Presence of trees >60>60 cm diameter at breast height in $\times4$ km cells (NFI based), indicating continuity and providing specific habitats.
@@
Nature Share 40% Landscape context where at least 40% of the surrounding $\times1$ km area is natural habitat, important for species connectivity.
Nature Share 80% Landscape context where at least 80% of the surrounding $\times1$ km area is natural habitat, indicating very high connectivity.
@@
Table 1: HNV forest proxy features. A short summary description of each of the 11 proxy features of the HNV forest map. we refer to (johannsen-hnv-skov-2015) for an exhaustive description.
In pilot experiments, we realized that it is very difficult to predict the exact HNV scores from remote sensing data. Thus, we simplified the task by distinguishing three aggregated classes, low, medium, and high biodiversity potential, each of which comprising roughly the same number of different HNV score values. We define low forest biodiversity potential as areas with HNV scores from 1–3, high forest biodiversity potential as areas with HNV scores from 7–10, and medium forest biodiversity potential as areas with HNV scores from 4–6. The focus of this study was on the binary task of distinguishing low and high biodiversity potential areas. Arguably, this is an important subproblem, as one wants to avoid mixing up these two classes in forest management decisions. Starting with simpler coarse-grained tasks that can be solved with an accuracy high enough to be regarded as informative and allows for a meaningful comparison of deep learning-based classification methods trained on 2D orthophotos, 3D ALS point clouds, and a combination of the two, while still evaluating the forest biodiversity assessment capabilities of deep learning-based methods.
@@
The 2D orthophotos used in this project are derived from yearly nationwide aerial photography of Denmark and are part of GeoDanmark’s geographic administration framework (geodanmarkforaarsbilleder). Orthophotos are created by adjusting each pixel according to the camera’s position and the elevation model of the Earth’s surface, ensuring that the images are geometrically corrected and can be used for precise measurements.
@@
The 3D ALS point clouds used in this study is part of the Danish Height Model dataset (dhmklimadatastyrelsen). The ALS product specifications can be found in the report by the dhm2020. This comprehensive dataset covers the entire country of Denmark and is continuously updated through yearly systematic airborne laser scanning. As of 2018 and forward, 1/5 of the country is updated annually. In this research we use ALS point clouds from 2019, 2020, 2021, 2022 and 2023 depending on the area of the forest, as illustrated in Figure 5. For each forest area, we ensure temporal consistency by using orthophotos captured in the same year as the corresponding ALS point cloud data, allowing for fair comparison between the two modalities and ensuring that both data sources represent the forest conditions from the same time period. The point clouds are available in tiles spanning $\times1$ km and the point density averages 8 points/m2 for data, with a horizontal accuracy of 0.15 m and a vertical accuracy of 0.05 m.